### v2

@@ F.2 HD 169142
Figure 17: Results of IPCA with ARDI (left) and rank-3 PCA with ADI-only (right), for an injected companion at a separation of $0\aas@@fstack{\prime\prime}31$ into the VLT/SPHERE data set of HD 169142 (Ren et al. 2023), using a contrast of $1.5\times 10^{-5}$. The position of the simulated candidate is indicated by a red square, while the location of the actual candidate proposed by Gratton et al. (2019); Hammond et al. (2023) is marked by a yellow square
Figure 18: Contrast curve at $5\sigma$ confidence for the HD 169142 data set, processed using PCA with ADI-only. The separation of the candidate at $0\aas@@fstack{\prime\prime}31$ is indicated by a black vertical line. The yellow star indicate the expected contrast of the candidate, inferred by Hammond et al. (2023).
@@ F.3 HD 97048
Figure 19: Results of IPCA (left), and rank-3 annular-PCA (right) with ADI-only, for a fake companion at a separation of $0\aas@@fstack{\prime\prime}45$, using a contrast of $3\times 10^{-5}$ injected into the VLT/SPHERE data set of HD 97048 (Ren et al. 2023). The location of the fake candidate is indicated by a red square, while the location of the actual protoplanet candidate inferred through the observation of kinks by Pinte et al. (2018) is indicated with a yellow square.
@@
Figure 20: Results of IPCA (left) and rank-2 PCA (right) both with ADI-only, for the uncropped observations of HD 163296. The position of the candidate, inferred by Izquierdo et al. (2022) and Pinte et al. (2018) from a local deviation in Keplerian velocity, is marked by the yellow square with dimensions $60\times 60$ pixels. A zoom into the $60\times 60$ pixels aperture is displayed in the bottom left of each image, with the planet’s location indicated by an arrow on the smaller images. Five simulated companions at a separation of $2\aas@@fstack{\prime\prime}2$ have been injected into the cube, with contrasts ranging from $1\times 10^{-3}$ to $3\times 10^{-6}$. White circle indicate their positions.
@@ F.4 HD 163296
Figure 21: Contrast curve at $5\sigma$ confidence for the HD 163296 data set, processed using PCA with ADI-only. The separation of the candidate, inferred by Izquierdo et al. (2022) and Pinte et al. (2018) from a local deviation in Keplerian velocity at $2\aas@@fstack{\prime\prime}2$, is indicated by a black vertical line.

### v3

@@ F.2 HD 169142
Figure 17: Results of IPCA with ARDI (left) and rank-3 PCA with ADI only (right), for an injected companion at a separation of $0\aas@@fstack{\prime\prime}31$ into the VLT/SPHERE dataset of HD 169142 (Ren et al. 2023), using a contrast of $1.5\times 10^{-5}$. The position of the simulated candidate is indicated by a red square, while the location of the actual candidate proposed by Gratton et al. (2019); Hammond et al. (2023) is marked by a yellow square
Figure 18: Contrast curve at $5\sigma$ confidence for the HD 169142 dataset, processed using PCA with ADI only. The separation of the candidate at $0\aas@@fstack{\prime\prime}31$ is indicated by a black vertical line. The yellow star indicate the expected contrast of the candidate, inferred by Hammond et al. (2023).
@@ F.3 HD 97048
Figure 19: Results of IPCA (left), and rank-3 annular-PCA (right) with ADI only, for a fake companion at a separation of $0\aas@@fstack{\prime\prime}45$, using a contrast of $3\times 10^{-5}$ injected into the VLT/SPHERE dataset of HD 97048 (Ren et al. 2023). The location of the fake candidate is indicated by a red square, while the location of the actual protoplanet candidate inferred through the observation of kinks by Pinte et al. (2018) is indicated with a yellow square.
@@
Figure 20: Results of IPCA (left) and rank-2 PCA (right) both with ADI only, for the uncropped observations of HD 163296. The position of the candidate, inferred by Izquierdo et al. (2022) and Pinte et al. (2018) from a local deviation in Keplerian velocity, is marked by the yellow square with dimensions $60\times 60$ pixels. A zoom into the $60\times 60$ pixels aperture is displayed in the bottom left of each image, with the planet’s location indicated by an arrow on the smaller images. Five simulated companions at a separation of $2\aas@@fstack{\prime\prime}2$ have been injected into the cube, with contrasts ranging from $1\times 10^{-3}$ to $3\times 10^{-6}$. White circle indicate their positions.
@@ F.4 HD 163296
Figure 21: Contrast curve at $5\sigma$ confidence for the HD 163296 dataset, processed using PCA with ADI only. The separation of the candidate, inferred by Izquierdo et al. (2022) and Pinte et al. (2018) from a local deviation in Keplerian velocity at $2\aas@@fstack{\prime\prime}2$, is indicated by a black vertical line.